The sum of the real numbers x and y is 11. Their difference is 5. what is the value of xy?
the answer is 24 but I do not understand how to get this please help!!! thank you
-Emalie

Answer by checkley75(3666)

X+Y=11 THEN X=11-Y & X-Y=5 NOW SUBSTITUTE (11-Y) FOR X IN THE SECOND EQUATION & SOLVE FOR Y:
(11-Y)-Y=5
11-Y-Y=5
-2Y=5-11
-2Y=-6
Y=-6/-2
Y=3 ANSWER.
X+3=11
X=11-3
X=8 ANSWER.
PROOF
3*8=24
24=24